OpenStudy (anonymous):

Expand the Logarithm Log6 2x/y

zepdrix (zepdrix):

The 6 is the base?

OpenStudy (anonymous):

yes :)

zepdrix (zepdrix):

We used our division rule in the last problem, now we want to use it in the reverse direction,\[\large\rm \log\left(\frac{a}{b}\right)=\log(a)-\log(b)\]

zepdrix (zepdrix):

Do you see how that will help us break the 2x away from the y?

OpenStudy (anonymous):

the 2x stays together?

zepdrix (zepdrix):

yes, for now.

OpenStudy (anonymous):

so Log (2x) - Log (y)??

zepdrix (zepdrix):

log(2x/y) = log(2x) - log(y) mmm k great, first step is done.

OpenStudy (anonymous):

or Log6 (2x) - Log6 (y)

zepdrix (zepdrix):

yes, we probably shouldn't ignore the base of the log :) my bad

OpenStudy (anonymous):

yeah haha :)

OpenStudy (anonymous):

then what?

zepdrix (zepdrix):

Then apply your multiplication rule to the 2 and x, log(ab) = log(a) + log(b)

OpenStudy (anonymous):

so log6 (2) + Log6 (x) ?

OpenStudy (anonymous):

?

zepdrix (zepdrix):

oh sorry got busy XD yes.

zepdrix (zepdrix):

log6(2x/y) = log6(2)+log6(x)-log6(y) yay team \c:/

OpenStudy (anonymous):

its okay :)

OpenStudy (anonymous):

is that it?